# Search for gravitationally lensed interstellar transmissions

### v1

@@ II.1 Light amplification
${\tt PSF}(\rho)$ $=2\pi kr_{g}J^{2}_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho\Big{)}\simeq 1.17\times 10^{11}J^{2}_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho\Big{)}\,\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}\Big{(}\frac{1\,\mu{\rm m}}{\lambda}\Big{)}.$ (9)
Fig. 3 shows the resulting Airy pattern of the transmitting lens projected on the solar system.
Note that an observer, positioned on the optical axis at $z\gg z_{0}$ from the gravitational lens, will see an Einstein ring around the lens with the radius $\theta_{\tt ER}$ given as:
$\theta_{\tt ER}=\frac{\sqrt{2r_{g}\tilde{z}}}{z}\simeq\frac{\sqrt{2r_{g}z_{0}}}{z}=2.46\times 10^{-9}\,\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{10\,{\rm pc}}{z}\Big{)}~{}{\rm rad}.$ (10)
@@
As a result, the spatial distribution characteristic to the PSF of the transmitting lens (9) and (13) will be encoded in the structure of the transmitted signal Turyshev (2024), guiding our search.
II.2 Power transmission
To assess the effectiveness of the interstellar power transmission with a stellar gravitational lens, we consider the same energy transmission scenario used to derive (1). We again assume that transmission is characterized by the beam divergence set by the telescope’s aperture $d_{\tt t}$ yielding angular resolution of $\theta_{0}\simeq\lambda/d_{\tt t}=1.00\times 10^{-6}~{}(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt t})\,{\rm rad}$. When the signal reaches the receiver at the distance of $(z_{\tt t}+z_{\tt r})$ from the transmitter, the beam is expanded to a large spot with the radius of $\rho_{*}=(z_{\tt t}+z_{\tt r})(\lambda/d_{\tt t})$. However, in this case, we need account for the fact that while passing by the lens, the light intensity is amplified according to (9). This is due to the fact that the lens collimates the otherwise diverging beam of light, thus delivering a larger power density to the receiver (see Fig. 1). As a result, a receiver telescope gets a larger fraction of the transmitted power
$P_{\tt R}^{\tt GL}$ $=$ $P_{\tt T}\frac{\pi({\textstyle\frac{1}{2}}d_{\tt r})^{2}}{\pi\rho_{*}^{2}}{\tt PSF}(\rho)=P_{\tt T}\frac{\pi({\textstyle\frac{1}{2}}d_{\tt r})^{2}}{\pi(z_{0}+z)^{2}}\Big{(}\frac{d_{\tt t}}{\lambda}\Big{)}^{2}\,2\pi kr_{g}J^{2}_{0}\big{(}k\theta_{\tt ER}\rho\big{)}\simeq$ (14)
$\simeq$ $3.06\times 10^{-13}J^{2}_{0}\big{(}k\theta_{\tt ER}\rho\big{)}\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}^{2}\Big{(}\frac{d_{\tt t}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{d_{\tt r}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}^{2}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}~{}{\rm W}.~{}$
Comparing (14) and (1), we see that power transmission link aided by a gravitational lens amplifies the received power by $P^{\tt GL}_{\tt R}/P^{\tt free}_{\tt R}\simeq 1.17\times 10^{11}J^{2}_{0}\big{(}k\theta_{\tt ER}\rho\big{)}$, as given by (9).
Amplification (×$\times 10^{10}$)
Distance from the optical axis, $\rho$ [m]
@@ II.2 Power transmission
Assuming that the receiver is near the optical axis, $\rho\simeq\rho_{\tt GL}$, to estimate the maximum photon flux at the receiver, we compute $Q_{\tt GL}=(\lambda/hc)P^{\tt GL}_{\tt R}$ from (14), yielding
$\hskip-2.0ptQ_{\tt GL}(\rho)\hskip-2.0pt$ $\simeq$ $\hskip-2.0pt1.54\times 10^{6}J^{2}_{0}\big{(}k\theta_{\tt ER}\rho\big{)}\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}\Big{(}\frac{d_{\tt t}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{d_{\tt r}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}^{2}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}~{}{\rm phot/s}.~{}~{}~{}~{}$ (15)
For large separations, $\rho\gg\rho_{\tt GL}$, we use (12) to transform (15) as below:
$Q_{\tt GL}(\rho)$ $\simeq$ $3.18\times 10^{4}\Big{(}\frac{1\,{\rm m}}{\rho}\Big{)}\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{d_{\tt t}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{d_{\tt r}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}~{}{\rm phot/s}.~{}~{}~{}~{}~{}$ (16)
@@
Therefore, thanks to gravitational lensing, there is a significant photon flux at the receiver. Now we need to consider the feasibility of such a transmission link. For this, we have to evaluate the contribution of major noise sources to the overall detection sensitivity.
II.3 Major noise source

### v3

@@ II.2 Light amplification
${\tt PSF}(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})$ $=2\pi kr_{g}J^{2}_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})\Big{)}\simeq 1.17\times 10^{11}J^{2}_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})\Big{)}\,\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}\Big{(}\frac{1\,\mu{\rm m}}{\lambda}\Big{)}.$ (18)
Fig. 3 shows the resulting Airy pattern formed by the lens in an image plane in the solar system. With this result, we establish the angular resolution of the transmitting lens which is determined by the first zero of the Bessel function $J_{0}(x)$ in (18) Turyshev and Toth (2020a), which occurs at $x=2.40483$, yielding the angular resolution measured in nano-arcseconds (nas):
$r_{\tt GL}=\big{|}\frac{{\boldsymbol{\mathrm{x}}}}{z}+\frac{{\boldsymbol{\mathrm{x}}^{\prime}}}{{z}_{0}}\big{|}=\frac{2.40483}{k\sqrt{2r_{g}\tilde{z}}}=0.38\frac{\lambda}{\sqrt{2r_{g}\tilde{z}}}\simeq 0.10\Big{(}\frac{\lambda}{1\,\mu{\rm m}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}^{\frac{1}{2}}~{}{\rm nas}.$ (19)
We observe that the form of resolution $r_{\tt GL}$ is similar to that of a classical diffraction-limited telescope that behaves as $\sim 1.22\,\lambda/d$. However, in the case of a gravitational lens, the denominator is the impact parameter $\sqrt{2r_{g}\tilde{z}}=b\gtrsim R_{\tt L}$, which is why result (19) yields a very high resolution behaving as $r_{\tt GL}\lesssim 0.38\,\lambda/R_{\tt L}$. This underscores the fact that after passing by the lens and traveling a large distance from it, the beam becomes tightly compressed.
An observer, positioned on the optical axis at $z\gg z_{0}$ from the gravitational lens, will see an Einstein ring around the lens with the radius $\theta_{\tt ER}$ given as: [^3] Note that in typical microlensing scenarios, where the source is at much larger distance $z_{0}$ from the lens then an observer is, $z$, the following is valid: $z_{0}\gg z$. In this case, the size of the Einstein ring is, $\theta_{\tt ER}={\sqrt{2r_{g}\tilde{z}}/{z}}\simeq\sqrt{{2r_{g}}/{z}}=1.38\times 10^{-7}\,\sqrt{M_{\tt L}/M_{\odot}}\,\sqrt{10~{}{\rm pc}/z}$ or $\sim{\cal O}(28.54~{}{\rm mas})$, which is much larger than that estimated in (20) and could be resolved by the largest instruments available Turyshev and Toth (2023).
$\theta_{\tt ER}=\frac{\sqrt{2r_{g}\tilde{z}}}{z}\simeq\frac{\sqrt{2r_{g}z_{0}}}{z}=2.46\times 10^{-9}\,\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{10\,{\rm pc}}{z}\Big{)}~{}{\rm rad},$ (20)
or $\sim{\cal O}(0.5~{}{\rm mas})$, which is quite challenging to resolve with the current generation of optical instruments (see Turyshev and Toth (2023)).
@@
Amplification (×$\times 10^{10}$) Refer to caption Distance from the optical axis, $\rho$ [m]
@@
As a result, the spatial structure of the transmitting lens’s PSF, as depicted in Fig. 3 and described by (18), will influence the received signal. This structure is well-characterized: at typical wavelengths, its spatial frequency in the image plane scales to hundreds of meters, as specified by (20), and its amplitude is inversely proportional to the radial distance from the center of the pattern, as illustrated by (23) and (32). These characteristics of the PSF’s spatial distribution (18) and (23) will be encoded in the structure of the received signal, thus directing our search strategy.
II.3 Power projection by an extended source
Examining the PSF (18) and recognizing that the angular resolution (19) is extremely small, we see that a monopole gravitational lens acts as a convex lens by focusing light but also by expanding the projected beam, according to ${\boldsymbol{\mathrm{x}}}=-({z}/{z_{0}}){\boldsymbol{\mathrm{x}}}^{\prime}$. This mapping rule suggests that by the time the transmitted beam reaches the solar system, its size will have increased by a factor of ${z}/{z_{0}}\sim 3.17\times 10^{3}\,({z}/10\,{\rm pc})(650\,{\rm AU}/z_{0})$. As a result, the diameter of the transmitting telescope as seen by an observer (i.e., the brightest spot) will be expanded to fill in a larger area of
$d_{\tt beam}=\frac{z}{z_{0}}d_{\tt T}\simeq 3.17\,\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}~{}{\rm km}.$ (24)
This is the diameter of the aperture of the transmitting telescope projected onto the image plane by the PSF of a monopole gravitational lens (18). It does not account for the diffraction of the transmitted beam; thus, it is much smaller compared to the diameter of the entire caustic given by (9). However, it gives the size of the beam with the highest intensity characteristic to the strong interference region.
Even the largest telescopes today will not be able to see the entire beam at once. (We dealt with a similar situation while considering the imaging of exoplanets with the SGL Turyshev and Toth (2020b).) Using (24), we see that a telescope with the aperture $d_{\tt R}\sim 10$ m could resolve the projected laser beam with diameter of $d_{\tt beam}$ with $N$ linear resolution elements given by
$N=\frac{d_{\tt beam}}{d_{\tt R}}=\frac{z}{z_{0}}\frac{d_{\tt T}}{d_{\tt R}}\simeq 317.33\,\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{d_{\tt T}}{1\,{\rm m}}\Big{)}\Big{(}\frac{10\,{\rm m}}{d_{\tt R}}\Big{)}.$ (25)
This behavior is consistent with the PSF broadening, as given by (18) and (21), and its impact on the expansion of a light beam emitted from the focal point of a convex lens exhibiting spherical aberration. With this insight, we need to treat a laser transmitter operating in the focal region of a gravitational lens as an extended source.
Therefore, we consider the laser transmitter to be an extended luminous source with surface brightness $B(\boldsymbol{\mathrm{x}}^{\prime})$ with dimensions of ${\rm W\,m}^{-2}{\rm sr}^{-1}$, the power density, $I_{0}(\boldsymbol{\mathrm{x}})$, received in the receiver plane is computed by integrating the PSF (18) over the surface of the extended source and is given as below (with $\theta_{\tt ER}$ from (20))
$I_{0}(\boldsymbol{\mathrm{x}})$ $=$ $\frac{2\pi kr_{g}}{(z_{0}+z)^{2}}\iint d^{2}\boldsymbol{\mathrm{x}}^{\prime}\,B(\boldsymbol{\mathrm{x}}^{\prime})\,J^{2}_{0}\Big{(}k\theta_{\tt ER}\big{|}{\boldsymbol{\mathrm{x}}}+\frac{z}{z_{0}}{\boldsymbol{\mathrm{x}}}^{\prime}\big{|}\Big{)}.$ (26)
A telescope with aperture $d_{\tt R}\ll d_{\tt sig}$, centered at a particular point $\boldsymbol{\mathrm{x}}_{0}$ in the image plane, will receive the signal $P(\boldsymbol{\mathrm{x}}_{0})=\iint d^{2}\boldsymbol{\mathrm{x}}\,I_{0}\big{(}\boldsymbol{\mathrm{x}}_{0}+\boldsymbol{\mathrm{x}}\big{)},$ where the integration is done within the telescope’s aperture $|\boldsymbol{\mathrm{x}}|\leq{\footnotesize\frac{1}{2}}d_{\tt R}$, and with $|{\boldsymbol{\mathrm{x}}}_{0}+{\boldsymbol{\mathrm{x}}}|\leq{\footnotesize\frac{1}{2}}d_{\tt beam}$. This yields a result that depends on the telescope’s position on the receiver plane:
$P_{\tt R}({\boldsymbol{\mathrm{x}}}_{0})$ $=$ $\frac{2\pi kr_{g}}{(z_{0}+z)^{2}}\hskip-3.0pt\iint\displaylimits_{|{\boldsymbol{\mathrm{x}}}|^{2}\leq(\frac{1}{2}d_{\tt R})^{2}}\hskip-3.0ptd^{2}\boldsymbol{\mathrm{x}}\hskip-5.0pt\iint\displaylimits_{|{\boldsymbol{\mathrm{x}}^{\prime}}|^{2}\leq(\frac{1}{2}d_{\tt T})^{2}}\hskip-5.0ptd^{2}\boldsymbol{\mathrm{x}}^{\prime}B(\boldsymbol{\mathrm{x}}^{\prime})\,J^{2}_{0}\Big{(}k\theta_{\tt ER}\big{|}{\boldsymbol{\mathrm{x}}}_{0}+{\boldsymbol{\mathrm{x}}}+\frac{z}{z_{0}}{\boldsymbol{\mathrm{x}}}^{\prime}\big{|}\Big{)}.$ (27)
In the case of a laser for which the emission is highly collimated which may conservatively be characterized by the beam divergence of $\theta_{0}=\beta(\lambda/d_{\tt T})$, with $\beta\simeq{\cal O}(1)$. As a result, the laser luminosity is scaled as $L_{0}=P_{\tt T}/\theta_{0}^{2}$. Although laser beams exhibit Gaussian spatial profile, for simplicity, we will assume a top-hat profile with uniform brightness, so, that the source’s surface brightness takes the form: $B(\boldsymbol{\mathrm{x}}^{\prime})=P_{\tt T}/\theta_{0}^{2}\pi(\frac{1}{2}d_{\tt T})^{2}$. Using this expression in (27), yields
$P_{\tt R}({\boldsymbol{\mathrm{x}}}_{0})$ $=$ $\frac{2\pi kr_{g}}{(z_{0}+z)^{2}}\frac{P_{\tt T}}{\pi(\frac{1}{2}d_{\tt T})^{2}}\Big{(}\frac{d_{\tt T}}{\lambda}\Big{)}^{2}\hskip-3.0pt\iint\displaylimits_{|{\boldsymbol{\mathrm{x}}}|^{2}\leq(\frac{1}{2}d_{\tt R})^{2}}\hskip-3.0ptd^{2}\boldsymbol{\mathrm{x}}\hskip-5.0pt\iint\displaylimits_{|{\boldsymbol{\mathrm{x}}^{\prime}}|^{2}\leq(\frac{1}{2}d_{\tt T})^{2}}\hskip-5.0ptd^{2}\boldsymbol{\mathrm{x}}^{\prime}\,J^{2}_{0}\Big{(}k\theta_{\tt ER}\big{|}{\boldsymbol{\mathrm{x}}}_{0}+{\boldsymbol{\mathrm{x}}}+\frac{z}{z_{0}}{\boldsymbol{\mathrm{x}}}^{\prime}\big{|}\Big{)},$ (28)
which, in the case when gravity is absent, is identical to (1). Otherwise, the difference is due to the PSF presence.
To evaluate this integral, we use the same approach developed in Turyshev and Toth (2020b, a). As a result, defining $\rho_{0}=|\boldsymbol{\mathrm{x}}_{0}|$, the power received at the detector placed in the focal plane of an optical telescope is given as
$P_{\tt R}(\rho_{0})$ $=$ $P_{\tt T}\frac{\pi({\textstyle\frac{1}{2}}d_{\tt R})^{2}}{\pi(z_{0}+z)^{2}}\Big{(}\frac{d_{\tt T}}{\lambda}\Big{)}^{2}\,\frac{4\sqrt{2r_{g}\tilde{z}}}{d_{\tt T}}\,\mu(\rho_{0}),\qquad{\rm with}\qquad\mu(r_{0})=\frac{2}{\pi}{\tt Re}\Big{\{}{\tt E}\Big{[}\Big{(}\frac{\rho_{0}}{{\textstyle\frac{1}{2}}d_{\tt beam}}\Big{)}^{2}\Big{]}\Big{\}},$ (29)
where, following (Turyshev and Toth, 2020a, b, c), we introduced the blur factor $\mu(\rho_{0})$ due to the spherical aberration of the monopole lens with ${\tt E}[x]$ being the elliptic integral (Abramowitz and Stegun, 1965) and ${\tt Re}\{\}$ is the operation taking the real part of a complex quantity.
The behavior of $\mu(\rho_{0})$ is shown in Fig. 6 of Turyshev and Toth (2020a). This factor captures blur’s contribution present on the image plane due to spherical aberration of a monopole lens. It reaches its maximal value of 1 for receiver on the optical axis, i.e., $\mu(0)=1$. On the edge of the projected beam, it decreases to $\mu({\textstyle\frac{1}{2}}d_{\tt beam})=0.64$. Outside the footprint of the beam projected on the solar system, $\mu(\rho_{0})$ falls off as $\propto{\textstyle\frac{1}{2}}d_{\tt beam}/\rho_{0}$, as expected from the PSF of a monopole lens Turyshev and Toth (2022a).
Comparing the results for a free-space propagation, given by (1), and that for gravitationally-lensed scenarios, from (29), we see that treating transmitter as an extended and resolved source, the light amplification is given as
${\cal A}_{\tt GL}(\rho_{0})$ $=$ $\frac{4\sqrt{2r_{g}\tilde{z}}}{d_{\tt T}}\,\mu(\rho_{0})=3.03\times 10^{9}\,\mu(\rho_{0})\,\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{1~{}{\rm m}}{d_{\tt T}}\Big{)},$ (30)
which is less compared to the amplification for point sources, given by (18), as expected, but still very high.
Eq. (29) describes the power of the light received from the transmitted laser beam that is present in the receiver plane, both inside and outside the beam’s footprint. Such behavior of the signal is due to the specific optical properties of the lens given by its PSF (18) which, as a function of the distance to the optical axis on the image plane, falls out much more slowly than the PSF of a regular telescope (see discussion in Turyshev and Toth (2020b)). This fact provides valuable insight for the search of a transmitted signal and the relevant work on prospective search campaign planning and development.
III Detection sensitivity
Understanding the detection sensitivity is crucial in the search for transmitted laser signals. This is done by evaluating the signal-to-noise ratio (SNR) in realistic transmission scenarios, which will be done below.
III.1 Evaluating anticipated signals
To assess the effectiveness of the interstellar power transmission with a stellar gravitational lens, we consider the same energy transmission scenario used to derive (1). We again assume that transmission is characterized by the beam divergence set by the telescope’s aperture $d_{\tt T}$ yielding angular resolution of $\theta_{0}\simeq\lambda/d_{\tt T}=1.00\times 10^{-6}~{}(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt T})\,{\rm rad}$. When the signal reaches the receiver at the distance of $(z_{\tt t}+z_{\tt r})$ from the transmitter, the beam is expanded to a large spot with the radius of $\rho_{*}=(z_{\tt t}+z_{\tt r})(\lambda/d_{\tt T})$. However, in this case, we need account for the fact that while passing by the lens, the light intensity is amplified according to (18). This is due to the fact that the lens collimates the otherwise diverging beam of light, thus delivering a larger power density to the receiver (see Fig. 1). As a result, based ion (29) a receiver telescope gets a larger fraction of the transmitted power
$P_{\tt R}^{\tt GL}$ $=$ $P_{\tt T}\frac{\pi({\textstyle\frac{1}{2}}d_{\tt R})^{2}}{\pi(z_{0}+z)^{2}}\Big{(}\frac{d_{\tt T}}{\lambda}\Big{)}^{2}\,\frac{4\sqrt{2r_{g}z_{0}}}{d_{\tt T}}\,\mu(\rho_{0})\simeq$ (31)
$\simeq$ $7.95\times 10^{-15}\,\mu(\rho_{0})\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}^{2}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{d_{\tt R}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}^{2}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}~{}{\rm W}.~{}$
Comparing (31) and (1), we see that power transmission link aided by a gravitational lens amplifies the received power by $P^{\tt GL}_{\tt R}/P^{\tt free}_{\tt R}\simeq 3.03\times 10^{9}\,\mu(\rho_{0})$, as given by (30).
@@ III.1 Evaluating anticipated signals
To estimate photon flux at the receiver, we compute $Q_{\tt GL}=(\lambda/hc)P^{\tt GL}_{\tt R}$ from (31), yielding
$\hskip-2.0ptQ_{\tt GL}(\rho)\hskip-2.0pt$ $\simeq$ $\hskip-2.0pt4.01\times 10^{4}\,\mu(\rho_{0})\Big{(}\frac{P_{\tt T}}{1~{}{\rm W}}\Big{)}\Big{(}\frac{1~{}\mu{\rm m}}{\lambda}\Big{)}^{2}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{d_{\tt R}}{1~{}{\rm m}}\Big{)}^{2}\Big{(}\frac{10~{}{\rm pc}}{z}\Big{)}^{2}\Big{(}\frac{M_{\tt L}}{M_{\odot}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}^{\frac{1}{2}}~{}~{}{\rm phot/s}.$ (32)
@@
Thus, due to gravitational lensing, the receiver experiences a significant photon flux. To assess the viability of this transmission link, it’s essential to evaluate how major noise sources affect the overall detection sensitivity.
III.2 Major noise source